# Open-Vocabulary Action Localization with Iterative Visual Prompting

### v4

@@ IV-D Action-order Aware Prompting
To estimate the start and end times of each task in parallel with Algorithm 2, each T-PIVOT process needs to know which task in the task sequence it is focusing on. Without this knowledge, the process cannot determine which instance to focus on; for example, if there are two instances of “picking up” in a video, it is not clear which “picking up” action is of interest when given a query of “Please find the picking up action in this video.” To address this issue, we prepared the following prompt using a sequence of task labels $L_{i}$ ($i=1\ldots N$):
@@
where {task_sequence} lists the sequence of tasks in the format “1. $L_{1}$, 2. $L_{2}$, …” and {task_focus} refers to the specific task being searched for in this instance (i.e., “k. $L_{k}$”). This prompt is designed to consider the order and focus of tasks during the process, making the VLM aware of which task is of the current interest.
@@ IV-E Performance Across Different Prompting Styles
Second, we investigated how to create images for inclusion in the prompt, specifically focusing on how many images from a time window should be presented and in what manner. Table II and Table III show the impact of grid size, which reflects the number of sampled images per iteration, on the results. The highest performance was achieved with a 5x5 grid (i.e., extracting 25 frames from a sampling time window) across the three metrics for the Breakfast Dataset, and 3x3- or 4x4-grid settings for the Fine-grained Breakfast Dataset. These results appear to highlight the trade-off between temporal and spatial resolution in this method—–a smaller grid size reduces the temporal resolution in recognition as it limits the number of frames that can be extracted from the video, while a larger grid size lowers the spatial resolution of each frame, assuming the number of iterations and the resolution of the prompt image remain the same.
Lastly, we investigated other parameters for visual prompting, including the rendering position of the annotation numbers and the tiling spacing (Fig. 4). We also explored an approach where the images were input into the model as stacked buffers rather than tiled. The highest performance across the three metrics was achieved using either the original tiling method or the stacking method. The inferior performance observed with center-rendering of the annotation might be due to overlapping annotated numbers reducing the image clarity.
@@
TABLE III: Performance Metrics Across Grid Sizes
(Fine-grained Breakfast Dataset)
@@
Furthermore, to gain deeper insights into the performance of action recognition, we examined the impact of varying numbers of action steps and durations of each video on performance (Fig. 5). These results indicate that the number of action steps significantly affects the outcomes, highlighting remaining challenges in recognizing actions in longer videos.
Figure 5: Action localization performance plotted against action step number and the video length. (a) Breakfast Dataset and (b) Fine-grained Breakfast Dataset.
@@ IV-F Comparison with Existing Methods
Table V compares our proposed method with existing methods on the Breakfast Dataset, with the reported numbers obtained from [36]. It is important to note that this comparison is not direct due to fundamental differences in experimental assumptions. T-PIVOT assumes knowledge of the action sequence in the video for task-transition estimation, while learning-based methods typically do not rely on this information, making them more challenging. Despite these differences, existing methods still outperform our approach, highlighting its limitations compared to state-of-the-art learning-based techniques.
To gain insights into the zero-shot capabilities of our method, we compared its performance with existing zero-shot action localization methods. To this end, we used the THUMOS14 dataset and its officially provided evaluation code [50]. Importantly, this task is not transition estimation because the THUMOS14 dataset does not assume that tasks are continuously present in the video. Thus, we employed a different approach from the one in subsection IV-C. We divided each video into fixed time intervals and applied Algorithm 1 within those intervals to localize the start and end of actions. Specifically, the video was divided into intervals of five seconds, and a 5x5 grid tiling with four iterations was applied. To operate T-PIVOT, we assumed that the action labels contained in the video were known. Following the methods in [14], we report the mean average precision (mAP), which represents the mean of the average precision values computed at different IoU thresholds ([0.3, 0.4, 0.5, 0.6, 0.7]). Table VI shows the comparison results, where the reported numbers were obtained from [14]. These results show that the proposed method achieves state-of-the-art performance. Importantly, the existing method employs pre-training even though the performance is tested under zero-shot conditions, while our method does not require any pre-training.
@@
The scores for the top and the second best performances are bolded and underlined, respectively.
@@
The scores for the top and the second best performances are bolded and underlined, respectively.
V Discussion and Limitation
In this paper, we propose a method for VLM-empowered OVAL, using GPT-4o as an example of VLMs. Despite its straightforward iterative pipeline, our method demonstrated reasonable performance in both third-person and first-person videos. The results showed reasonable performance in task-transition estimation, with limitations such as decreased performance on videos with longer action sequences. Furthermore, the method demonstrated performance comparable to state-of-the-art zero-shot action localization. Importantly, the unique advantage of the proposed method lies in its training-free nature, eliminating the need for data collection or training. This allows for immediate and effortless application to new vocabularies and dataset domains, offering adaptability to diverse applications such as video annotation and editing.
The comparison of different prompting styles has revealed several limitations of this method. First, the optimal number of frames for sampling depends on a trade-off between spatial and temporal resolution in action recognition. This number may vary based on the video’s length and the nature of the actions, potentially requiring adjustments for each use case. Second, the method’s performance is negatively impacted by the length of action sequences and by numerical annotations that interfere with image clarity. These issues highlight opportunities to improve the method by adjusting the placement of number renderings or by modifying the video length beforehand, as employed in the zero-shot experiment. Lastly, since our method requires multiple queries to the VLM, it takes a longer time to recognize actions compared to typical learning-based methods. This issue could potentially be mitigated by using a more lightweight or local model that does not rely on network communication.
@@ VI Conclusion
This paper introduces a novel approach for open-vocabulary video action localization leveraging the PIVOT methodology, which eliminates the need for data collection and model training. Testing with benchmark datasets has demonstrated that the proposed method offers promising zero-shot capabilities without the need for model pre-training. Considering several limitations of this method, future work will focus on refining visual prompting techniques and optimizing the system to enhance its applicability across a broader range of video analysis tasks, including automated video labeling and advanced robotics training.

### v5

@@ IV-D Action-order Aware Prompting
To estimate the start and end times of each task in parallel with Algorithm 2, each T-PIVOT process needs to know which task in the task sequence it is focusing on. Without this knowledge, the process cannot determine which instance to focus on; for example, if there are two instances of “picking up” in a video, it is not clear which “picking up” action is of interest when given a query of “Please find the picking up action in this video.” To address this issue, we designed the following prompt using a sequence of task labels $L_{i}$ ($i=1\ldots N$):
@@
Here, {task_sequence} lists the sequence of tasks in the format “1. $L_{1}$, 2. $L_{2}$, …” and {task_focus} refers to the specific task being searched for in this instance (i.e., “k. $L_{k}$”). This prompt is designed to make the VLM aware of both the task order and which task is currently being targeted.
@@ IV-E Performance Across Different Prompting Styles
Second, we investigated how to create images for inclusion in the prompt, specifically focusing on how many images from a time window should be presented and in what manner. Table II and Table III show the impact of grid size, which reflects the number of sampled images per iteration, on the results. The highest performance was achieved with a 5x5 grid (i.e., extracting 25 frames from a sampling time window) across the three metrics for the Breakfast Dataset, and 3x3- or 4x4-grid settings for the Fine-grained Breakfast Dataset. These results appear to highlight the trade-off between temporal and spatial resolution in this method—a smaller grid size reduces the temporal resolution in recognition as it limits the number of frames that can be extracted from the video, while a larger grid size lowers the spatial resolution of each frame, under the assumption that the number of iterations and the resolution of the prompt image remain the same.
Lastly, we investigated other parameters for visual prompting, including the rendering position of the annotation numbers and the tiling spacing (Fig. 4). We also explored an approach where the images were input into the model as stacked buffers rather than tiled. The highest performance across the three metrics was achieved using either the original tiling method or the stacking method. The inferior performance observed with the center rendering of the annotation numbers might be due to the rendered numbers obscuring important visual content, making it harder to interpret the underlying image.
@@
TABLE III: Performance Metrics Across Grid Sizes (Fine-grained Breakfast Dataset)
@@
Furthermore, to gain deeper insights into the performance of action recognition, we examined the impact of the number of actions in each video and the duration of each video on performance (Fig. 5). These results indicate that the number of action steps significantly affects the outcomes, highlighting the remaining challenges in recognizing actions in longer videos.
Figure 5: Action localization performance plotted against the number of action steps and the video length. (a) Breakfast Dataset and (b) Fine-grained Breakfast Dataset.
@@ IV-F Comparison with Existing Methods
Table V compares our proposed method with existing methods on the Breakfast Dataset, with the reported numbers obtained from [36]. It is important to note that this comparison is not direct due to fundamental differences in experimental assumptions. T-PIVOT assumes the action sequence in the video is given for task-transition estimation. In contrast, learning-based methods typically operate without this information, making the problem setting more challenging for them. Despite these differences, existing methods still outperform our approach, highlighting its limitations compared to the state-of-the-art learning-based techniques.
To gain insights into the zero-shot capabilities of our method, we compared its performance with existing zero-shot action localization methods. To this end, we used the THUMOS14 Dataset and its officially provided evaluation code [53]. Importantly, this task is not transition estimation because the THUMOS14 Dataset does not assume that tasks are continuously present in the video. Thus, we employed a different approach from the one in Sec. IV-C. Similar to Sec. III-C, we divided each video into fixed time intervals and applied Algorithm 1 to each interval to localize the start and end of actions. Specifically, the video was divided into intervals of five seconds, and a 5x5 grid tiling with four iterations was applied. To operate T-PIVOT, we assumed that the action labels contained in the video were known. Following the methods in [14], we report the mean average precision (mAP), which represents the mean of the average precision values computed at different IoU thresholds (i.e., [0.3, 0.4, 0.5, 0.6, 0.7]). Table VI shows the comparison results, where the reported numbers were obtained from [14]. These results show that the proposed method achieves state-of-the-art performance. Importantly, the existing method employs pre-training even though the performance is tested under zero-shot conditions, while our method does not require any pre-training.
@@
The top and second-best scores are bolded and underlined, respectively.
@@
The top and second-best scores are bolded and underlined, respectively.
V Discussion and Limitations
In this paper, we proposed a method for VLM-empowered OVAL, using GPT-4o as an example of VLMs. Despite its straightforward iterative pipeline, our method demonstrated reasonable performance in both third-person and first-person videos. The results were promising in task-transition estimation, although limitations remain—for example, performance tends to degrade on videos with longer action sequences. Additionally, the proposed method showed competitive results among existing zero-shot methods. Notably, the unique advantage of the proposed method lies in its training-free nature—while the prior methods rely on pre-training even under zero-shot conditions, our method requires no model training nor data collection, allowing for immediate and flexible application to new vocabularies and dataset domains. Such adaptability makes it suitable for various applications such as video annotation and editing.
V-A Limitation
The T-PIVOT method involves several hyperparameters in visual prompting style, including language prompts, grid size, and the number of iterations, which may require tuning for each use case. Nevertheless, the experiments on different prompting styles in Sec. IV-E have revealed the general trends. First, the iteration number tends to improve performance up to a certain point. Next, the optimal number of frames for sampling depends on a trade-off between spatial and temporal resolution—a smaller grid size reduces temporal resolution by limiting the number of extracted frames, whereas a larger grid size lowers spatial resolution per frame, under the assumption that the number of iterations and the prompt image resolution remain the same. Finally, the method’s performance is impacted by the rendering position of the annotation numbers, suggesting the importance of rendering that avoids interference with image clarity. These results provide practical guidance for designing visual prompting.
As another limitation, the performance of T-PIVOT can also be influenced by the choice of VLM. To further examine its impact, we conducted an additional experiment using Phi-3-vision [64], one of the state-of-the-art open-source VLMs, applying it to the Breakfast Dataset under the same conditions as GPT-4o. Phi-3-vision showed considerably lower performance (MoF: 31.9, IoU: 18.5, and F1: 25.4) compared to GPT-4o (MoF: 60.1, IoU: 40.8, and F1: 51.2; see Table V). While open-source VLMs generally perform well in describing single images [65], T-PIVOT requires extracting temporal information from a single image that summarizes an image sequence. Furthermore, this task demands interpretation of textual prompts to identify the start and end points of a specified action, making it more complex than standard image description. These results indicate that our approach benefits from models with strong video reasoning capabilities.
Additionally, since our method requires multiple queries to the VLM, it tends to take more time compared to typical learning-based methods. This increased latency is a trade-off for the flexibility and adaptability provided by open-vocabulary models, which do not require task-specific training. While this drawback may limit some real-time applications, our method remains feasible for offline analysis or tasks where flexibility in specifying target actions is prioritized.
While our method involves hyperparameter tuning and relies on VLM performance, its novelty lies in the use of off-the-shelf VLMs for OVAL. Our primary contribution is bridging the multimodal perception capabilities of pre-trained VLMs with action localization through an iterative prompting strategy.
@@ VI Conclusion
This paper introduced a novel approach for OVAL leveraging the PIVOT methodology, which eliminates the need for data collection and model training. Testing with benchmark datasets has demonstrated that the proposed method offers promising zero-shot capabilities without the need for model training. Considering several limitations of this method, future work will focus on refining visual prompting techniques and optimizing the system to enhance its applicability across a broader range of video analysis tasks, including automated video labeling and advanced robotics training.
Acknowledgment
We thank Dr. Sakiko Yamamoto, Dr. Etsuko Saito (Ochanomizu University), and Dr. Midori Otake (Tokyo Gakugei University) for their help in annotating the Fine-grained Breakfast Dataset. This study was conceptualized, conducted, and written by the authors, and an AI, OpenAI’s GPT-4o, was used for proofreading.